OpenStudy (anonymous):

@cody7744 MEDAL AND FAN Complete the standard form of the equation 3y/2 = 8 - x Answer: ?x + ?y = 16 ( FILL IN )

OpenStudy (cj49):

3y/2=8-x 3y=8-2x 2x+3y=8 (to make 8=16 multiply the whole equation by 2) 2(2x+3y=8)

OpenStudy (anonymous):

ARE YOU 100% SURE

OpenStudy (cj49):

yes

OpenStudy (anonymous):

has to = 16

OpenStudy (anonymous):

?x + ?y = 16

OpenStudy (anonymous):

Yes he is right. Just multiply 2 on both sides. 4x+6y = 16

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

What is the x-intercept of the line with the equation 8x - 1/3y = 16 ( ? , 0 )

OpenStudy (cj49):

substitute y=0 in the equation 8x=16 x=16/8=?

OpenStudy (anonymous):

( ? , 0 )